A butcher has two grades of ground meat. A low grade meat costs $1.50 per pound and a higher grade of meat which costs $4.00 per pound. How many pounds of each grade of meat should he mix together to get 25 pounds of meat costing $3.00 per pound?
Answer by checkley77(12844)

4X+1.5(25-X)=3*25
4X+37.5-1.5X=75
2.5X=75-37.5
2.5X=37.5
X=37.5/2.5
X=15 POUNDS OF $4 MEAT IS USED.
25-15=10 POUNDS OF $1.50 MEAT IS USED.
PROOF:
4*15+1.5*10=3*75
60+15=75
75=75